OpenStudy (anonymous):

an equal arm balance is leveled when 20 washers are on 1 side and 10 bolts on the other. 4 bolts are added to 1 side. How many washers must be added to the other side? I just need to know the formula..thanks!

hero (hero):

Basically, this is the set up \[\frac{20 \space \text{washers}}{x \space \text{washers}} = \frac{10 \space \text{bolts}}{4 \space \text{bolts}}\]

hero (hero):

You can consider both sides of the equal sign to be opposite sides

hero (hero):

You can cancel washers and bolts to simply get: \[\frac{20}{x} = \frac{10}{4}\] You can simplify 10/4 to get \[\frac{20}{x} = \frac{5}{2}\] Then cross multiply to get 40 = 5x And finally divide both sides to get 40/5 = x 8 = x Therefore, 8 washers must be added to the other side

OpenStudy (anonymous):

thank you!

hero (hero):

yw